\underline{Quiz 4}

Davis Name: M212 Pledge:

(9pts.) 1. Use integration tables to integrate the following:

a. ∫ \frac{y^{2}+4y+6}{y^{2}+4y+5}dy b. ∫ (x^{3} − 5x + 2)e^{−2x}dx c. ∫ sin^{4} (x) cos^{2} (x)dx

(11pts.) 2. Match the slope fields with their differential equations. Then choose two of the differential equations to solve algebraically by separation of variables. a. \frac{dy}{dx} = y + x^{2}y b. \frac{dy}{dx} = xy^{2} sin (x^{2}) c. \frac{dy}{dx} = y + xy d. \frac{dy}{dx} = xe^{y}